OpenStudy (anonymous):

Make b the subject. 1/f = 1/a + 1/b

OpenStudy (akshay_budhkar):

1/f-1/a=1/b a-f/af=1/b therefore b=af/a-f

OpenStudy (anonymous):

okay I get the first step. Not the second.

OpenStudy (akshay_budhkar):